Spotlight
Facilities
Head
Faculty
Coordinators and Dy-Coordinators
User Committees
Staff
Student Homepages
How to Reach Us
IITR Home
Instrument Status
Log Reports
Status
S.No. Equipment Current Status Remarks Contact Details (STD Code: 01332)
1 FE-SEM (FEI) Working EDS Not Working Mr. Amit K. Sharma (284922)
2 TEM Working   Mr. Amit K. Sharma (284922)
3.  XRF Working   Mr. Ravindra Saini  (286272)
4.  EPMA Working   Mr. Ravindra Saini  (286272)
5. Powder XRD (Bruker) Working   Mr. Shiv Kumar (286302)
6. Thin-Film XRD (Regaku) Working   Mr. Mohd. Danish (286302)
7.  Single Crystal XRD (Bruker) Working   Prof. U.P. Singh (285329)
8. ICPMS (Old) Working   Mr. Kamal S. Gotyan (285384)
9. FLS Working   Mr. Kamal S. Gotyan (285384)
10.  ICPMS (New) Working Newly Installed Feb 2020 Mr. Kamal S. Gotyan (285384)
11.  SEM (Old) Not Working Mag: upto 10000X Mr. Amit K. Sharma (284922)
12. VSM (Old) Working   Dr. Vipin Chawla (284580)
13. DTA-TGA (3 Nos.) Working   Mr. Anil K. Saini (286293)
14. FE-SEM (Carl-Zeiss) Working EDS is working Mr. Amit K. Sharma (284922)
15. FE-SEM (Carl-Zeiss New) Working EDS is working Mr. Amit K. Sharma (284922)
16.  XPS Working   Mr. D. Prasad (286267)
17. VSM (New) Working   Dr. Vipin Chawla (284580)
18. PPMS Working   Dr. Vipin Chawla (284580)
19. AFM Working   Mr. D. Prasad (286267)
20. SQUID Not Working Replacement is under process April 2020 Mr. D. Prasad (286267)
21. NMR Working   Dr. Jay S. Meena (286170)
22. MCU Working   Prof. Pravindra Kumar (286312)
23. EPR Not Working   Dr. Jay S. Meena (286170)
24. LN2 Plant Working Another LN2 plant under installation Mr. Kamal S. Gotyan (285384)
25. MP-AES Working Newly Installed Jan 2020 Mr. Kamal S. Gotyan (285384)